OpenStudy (highschoolmom2010):

Suppose that the math team at your school competes in a regional tournament. The math team has 12 members. Regional teams are made up of 4 people. How many different regional teams are possible?

OpenStudy (kropot72):

You need to find the number of combinations of 12 people taken 4 at a time which is: \[12C4=\frac{12!}{4! \times 8!}=\frac{12\times11\times10\times9}{4\times3\times2}=you\ can\ calculate\]

OpenStudy (highschoolmom2010):

yea i was unsure of which i needed to use....

OpenStudy (highschoolmom2010):

thanks @kropot72

OpenStudy (austinl):

You probably have an \(\large{_nC_r}\) button on your calculator :)

OpenStudy (highschoolmom2010):

i dont have a calculator that has that i have the standard calculator

OpenStudy (kropot72):

You're welcome :)

OpenStudy (highschoolmom2010):

\[\frac{ 11880 }{ 24 }=495\]

OpenStudy (austinl):

Yep :)

OpenStudy (kropot72):

Your answer is correct. Good work!